OpenStudy (anonymous):

8 1/5 + 3 5/8 its just adding mixed numbers

OpenStudy (zepp):

Can you first change it into improper fractions?

OpenStudy (zepp):

E.g.: \(\large c\frac{a}{b} = \frac{cb+a}{b}\)

OpenStudy (anonymous):

11 33/40

OpenStudy (zepp):

8 1/5 would be 8*5+1 / 5 = 41/5, in improper fraction form 3 5/8 would be 3*8+5 / 8 = 29/8 Now what you need to do is to get the same denominator It would be 40 in this case, since it's the LCD \(\huge \frac{41*8}{5*8}+\frac{29*5}{8*5}\)